OpenStudy (anonymous):

A restaurant has one type of milk with 2% fat and another has 7% fat. How many quarts of each type does the restaurant to make 40 quarts of a milk mixture that is 5% fat? 24 quarts of the 2% milk and 16 quarts of the 7% milk 16 quarts of the 2% milk and 24 quarts of the 7% milk 16 quarts of the 2% milk and 16 quarts of the 7% milk 24 quarts of the 2% milk and 24 quarts of the 7% milk

OpenStudy (anonymous):

Milk= x Fat = 40-x 0.02 x + 0.07(40 -x ) =40(0.05) 2x + 280-7x = 200 -5x =-80 x = 16 fat = 40-16 = 24 Answer B, 16 quarts of the 2% milk and 24 quarts of the 7% milk